OpenStudy (anonymous):

how would i convert r = 1 + 2 cos(2θ) into the rectangular coordinates?

ganeshie8 (ganeshie8):

\(\large r = 1+ 2\cos (2\theta) = 1 + 2[2\cos^2\theta - 1] = 4\cos^2\theta - 1\) \(\large \implies r^3 = 4r^2\cos^2\theta - r^2 \)

ganeshie8 (ganeshie8):

plugin \(x = r\cos \theta \) \(r^2 = x^2+y^2\)

OpenStudy (anonymous):

does that mean r^3=4x^2 -x^2 +y^2?

ganeshie8 (ganeshie8):

Yep ! replace the \(r^3\) on left hand side also

OpenStudy (anonymous):

how would i do that?

ganeshie8 (ganeshie8):

\(\large r^3 = 3x^2 - y^2 \)

ganeshie8 (ganeshie8):

square both sides

ganeshie8 (ganeshie8):

\(\large (r^3)^2 = ( 3x^2 - y^2)^2 \)

ganeshie8 (ganeshie8):

\(\large (r^2)^3 = ( 3x^2 - y^2)^2 \)

ganeshie8 (ganeshie8):

\(\large (x^2+y^2)^3 = ( 3x^2 - y^2)^2 \)

ganeshie8 (ganeshie8):

we're done.

OpenStudy (anonymous):

but how would i do the graph for that equation?

ganeshie8 (ganeshie8):

graphing may be easier in polar coordinates itself : http://www.wolframalpha.com/input/?i=plot+r%3D4cos%5E2%28%5Ctheta%29-1

ganeshie8 (ganeshie8):

\(r = 1 + 2 \cos(2\theta )\) when \(\theta = 0^{\circ}\), \(r = 1+2(1) = 3\) so the curve starts on right side at r = 3

ganeshie8 (ganeshie8):

may be, plot few more points till u get some feel of how it is looping

OpenStudy (anonymous):

oh right! thanks

ganeshie8 (ganeshie8):

goodluck ! @BSwan is expert in sketching polar curves, Bswan give her a hand if you have time :)

OpenStudy (anonymous):

she wanna draw it ?

ganeshie8 (ganeshie8):

ask her

ganeshie8 (ganeshie8):

@swift_13

OpenStudy (anonymous):

oh its ohk i found out how to do it,thanks though

OpenStudy (anonymous):

:D good luck !

OpenStudy (anonymous):

even thought i love this curve xD